How much heat must be absorbed by 375 grams of water to raise its temperature by 25 degrees celsius? The specific heat of water is 4.186 J/ g degrees Celsius
1
Expert's answer
2019-05-07T08:09:55-0400

"Q=mc \Delta T"

"Q" - heat content

"m" - mass

"c" - specific heat

"\Delta T" - temperature change

"Q = 375 \cdot 4.186 \cdot 25 \approx 3.92 \cdot 10^4 \; J = 39.2 \; kJ"

Answer: 39.2 kJ.